OpenStudy (anonymous):

Write the equation of a line which has a slope of -4 and passes through the point (-5,1) ? What is the answer ? Thank you!!!

OpenStudy (anonymous):

m= -4 You'll have to consider the equation y=mx+b. then you simply plug in the points and the slope. 1 = -4*-5 + b 1 = 20 +b b = -19 y = -4x -19